ohyhyh:

The table below compares the radioactive decay rates of two materials. Material Original mass of material (in grams) Mass of material after 60 hours (in grams) 1 208 13 2 200 50 Based on the table, which of these conclusions is most likely correct? The half-life of Material 1 and Material 2 are equal. The half-life of Material 2 is double the half-life of Material 1. The half-life of Material 2 is 30 hours more than the half-life of Material 1. The half-life of Material 1 is 30 hours more than the half-life of Material 2.

ohyhyh:

im taking a FIFTY question assignment btw lol. Lot of tuff things and i need to have a passing grade.

KyanTheDoodle:

This is homework?

ohyhyh:

yes.

ohyhyh:

my teachers give a lot

ohyhyh:

im just putting anything that i cant get on here.

ohyhyh:

want to understand and get it correct

KyanTheDoodle:

Okay. Try using this formula: \[N(t)=N _{0}(\frac{ 1 }{ 2 })^{\frac{ t }{ t _{1/2} }}\] Where N(t) is the amount now present, NO is the initial amount, t is the time elapsed and t1/2 is the halflife.

KyanTheDoodle:

You know everything variable in this except for the half lifes. Try plugging in the information for both samples and see if there's anything different about them.

KyanTheDoodle:

*every variable

ohyhyh:

what do you mean byamount elapsed

ohyhyh:

by amount*

ohyhyh:

sorry btw. my 2 worst subjects are math and science haha

KyanTheDoodle:

so right next to the original amount, there's another amount. that's how much is left after the time passed.

KyanTheDoodle:

nah, you're good.

ohyhyh:

the thing i dont get it would the half like just be NO divided by 2

KyanTheDoodle:

that would be the amount left after its half life. the half life is how long it takes to get to that point.

KyanTheDoodle:

the amounts in this table undergo multiple half lifes so it's hard to calculate without a formula.

ohyhyh:

im confused. I got 3.25 for the 208

mhanifa:

You can calculate the half life in steps if the formula looks complicated. 1. Divide the initial amount by final one. 2. Identify it as powers of 2. 3. Divide 60 hours by the amount of power.

ohyhyh:

Identify "it" (?) as the powers of 2?

mhanifa:

For the first one you get: 208:13 = 16 16 = 2^4 60/4 = ...

ohyhyh:

15

mhanifa:

Same for the second one...

ohyhyh:

that would mean the second one is what 60? or 30?

mhanifa:

200/50 = 4 4 = 2^2 60/2 = ...

ohyhyh:

30

ohyhyh:

which would be b

mhanifa:

Yes , now you can read the options and see which is correct

ohyhyh:

thx

ohyhyh:

so who doi i best response lol

mhanifa:

@ohyhyh wrote:
which would be b
Correct

mhanifa:

@ohyhyh wrote:
thx
You are welcome!

ohyhyh:

thanks for helping guys both of u